OpenStudy (anonymous):

Use a double-angle formula to solve the following for x: cos 2x + cos x = 0

OpenStudy (anonymous):

recall that \[\cos(2x)=1-2\cos^2(x)\] so you get a quadratic equation in cosine

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

like solving \[1-2z^2+z=0\] you good from there?